OpenStudy (anonymous):

If LM=6 and HI=36, calculate the length of diameter LO answers a 54 b 18 c 30 d 60

OpenStudy (kenljw):

If you make ON = LM then you have a right triangle HIJ LO^2 = HI^2 + (LO - 2 LM)^2 LO^2 = HI^2 + LO^2 - 2 LO LM + 4 LM ^2 LO = (HI^2 + 4 LM^2)/2 LM = 120 Doesn't match any answer, not sure what's wrong